Charlie is 15 years older than Brett. Diane is 46. Paco is 10 years younger than Diane and 33 years older than Brett. How old is Charlie?

Answer by 1335(25)

Let's say Charlie is c, Diane is d, Paco is p, and Brett is b. Now plug what you know for each value: c=b+15; d=46; p=d-10; p=b+33.
First solve for Paco's age (p=d-10), where d=46: (p=46-10) = p=36.
Now use Paco's age to solve for Brett's age (p=b+33), where p=36
(36=b+33) = b=3.
Now use Brett's age to find Charlie's age: (c=b+15), where b=3
(c=3+15) = c=18
Charlie is 18 years old.